1. From the give table calculate Elasticity of Price, Total Revenue and Marginal Revenue. 

Also, explain the relationship between AR and MR?

6 0

5 100

4 200

3 300

2 400

1 500

0 600

Price(6-0), Quantity demand (0-600)


1
Expert's answer
2021-05-17T10:38:22-0400

price elasticity of demand

"Ed\u200b=\\frac{\\%\u0394P}{\\%\u0394Q}\u200b=\\frac{\\frac{Q_2-Q_1}{0.5(Q_2+Q_1)}}{\\frac{P_2-P_1}{0.5(P_2+P_1)}}"


"E_d\u200b=\\frac{\\frac{100-500}{0.5(100+500)}}{\\frac{5-1}{0.5(5+1)}}\u200b\u200b=\u22121.0"


the total revenue"=P\times Q"

"TR_1=6\times 0=0"

"TR_2=5\times100=500"

"TR_3=4\times 200=800"

"TR_4=3\times 300=900"

"TR_5=2\times400=800"

"TR_6=1\times 500=500"

"TR_7=0\times 600=0"


 the marginal revenue"=\frac {\Delta TR }{\Delta Q}"

"MR_1=\frac{0-500}{0-100}=5"


"MR_2=\frac{500-800}{100-200}=3"


"MR_3=\frac{800-900}{200-300}=1"


"MR_4=\frac{900-800}{300-400}=-1"


"MR_5=\frac{800-500}{400-500}=-3"


"MR_6=\frac{500-0}{500-600}=-5"




The marginal revenue is lower than the average revenue. Given the demand for this product, the monopolist can increase the sales by lowering the price and the marginal revenue also fall.